OpenStudy (anonymous):

I'm trying to find the limit as x approaches 4, abs(x-5) / x-5

hartnn (hartnn):

just plug in x = 4 in your function you know how to evaluate abs (4-5) ... ?

OpenStudy (anonymous):

abs(4-5) = abs(-1) = 1 4-5 = -1 1/(-1) = -1 ASNWER

OpenStudy (anonymous):

Sorry, I meant as the limit is approaching 5

hartnn (hartnn):

so you know what are left hand and right hand limits ?